OpenStudy (anonymous):

29. A recipe for 24 cookies calls for 6 tablespoons of sugar. If you make 36 cookies and use 10 tablespoons, will the cookies taste the same? Explain (2 points)

OpenStudy (anonymous):

do you do connections?

OpenStudy (anonymous):

use ratios : 6/24 = 1/4. the other recipe thingy is 10/36= 5/18. thus they will taste different because 5/18 is not equal to 1/4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i do